alones (alones):

3x + 5y = 78 2x - y = 0. I NEED ;[

OpenStudy (anonymous):

I figure you might solve that bottom equation for either y or x and then put that value into one of the spots in the top equation. Like, the bottom equation would be y = 2x or x = y/2. After that, I'm pretty sure you plug in one of those in place of x or y and solve for the one you are missing. Not entirely sure, but that's what it looks like to me.

OpenStudy (toloaf):

@Oris is right y=2x 3x+5(2x)=78 <---- plugin it in 13x=78 x=78/13=6 x=6 y=2(6) <---- plugin it in y=12 There you have it. x=6 and y=12 :)

alones (alones):

Thbaks you guys<33

OpenStudy (toloaf):

No problem! Glad we could help! ^^

alones (alones):

yea i love the way u explained thnaks :D

OpenStudy (toloaf):

Haha np Thanks for the testimonial =)

alones (alones):

YOUR MORE THANK WELCOMED =]